OpenStudy (anonymous):

2 grams of Magnesium is heated producing 3.33 grams of magnesium oxide. calculate the moles of magnesium present

OpenStudy (anonymous):

\[3.33 g MgO \left(\begin{matrix}1 mol MgO \\ 40.3 g MgO\end{matrix}\right)\left(\begin{matrix}24.3 g Mg \\ 1 mol MgO\end{matrix}\right)\left(\begin{matrix}1 mol Mg \\ 1 mol MgO\end{matrix}\right) = 2.00 mol Mg\]